# Feature-based Low-Rank Compression of Large Language Models via Bayesian Optimization

### v1

@@ 5.3 The Transferability of Rank Allocation
5.4 The Generalization of Validation Data
Wikitext ($\downarrow$) PTB ($\downarrow$) Zero-shot ($\uparrow$)
Wikitext 7.61 48.37 62.14
C4 7.65 44.56 62.07
Arxiv 8.46 46.77 62.11
Wikipedia 7.74 45.54 61.93
Table 4: Results under different validation data.
@@ Appendix A More Experiments on Low-rank Sensitivity
Appendix B Implementation Details
@@
At the post-training stage, we only add fine-tunable low-rank matrices for the compressed parameters. We set the low-rank dimension $r^{\prime}=256$, the learning rate is 1e-2, and the batch size is 64.
Appendix C Discussion on compute intensive about Bolaco
@@
PCA decomposition
Our method requires only one PCA decomposition of the obtained representations and truncates them according to the assigned rank during the rank allocation process to generate various compressed models. The computational cost here is the same as that of the existing low-rank decomposition method LORD.
Obtaining evaluation results
In our experiments, the validation set we selected is not large, about 34k tokens, so the validation process takes less time, and the total validation time spent by llama-2-7b is about 40-45min on a 40G A100.
Bayesian Optimization
The computational time for 50 epochs of Bayesian optimization is approximately 45-50 minutes, which is considered acceptable in practical applications. Compared to iterative pruning, Bayesian optimization is more memory-efficient, as it only requires the memory overhead of forward propagation without storing gradients, momentum, or other optimizer states. Furthermore, as discovered in Section 5.5, the low-rank configurations optimized on a base model can be transferred directly, or with few rounds of Bayesian optimization, to variant models with the same architecture. It implies that we can quickly obtain a well-performing, low-rank compressed model for fine-tuned LLMs on different datasets in practice.
Appendix D Statistics of the Compressed Model
@@
LLaMA-v2-7b Bolaco ($5\times 1$) 20% [824, 1696, 2208, NA, 2592]
Bolaco ($5\times 4$) 20% [[992, 1640, NA, NA, NA],
[824, 1808, 2360, NA, 2352],
[976, NA, 2696, 2400, 2736],
@@
LLaMA-v2-13b Bolaco ($5\times 1$) 20% [632, 1400, 2184, NA, 2976]
Bolaco ($5\times 4$) 20% [[816, 1536, 2512, 2984, NA],
[408, 1496, 2696, NA, 2960],
@@
Appendix E Language Modeling Capabilities for Compressed Models
@@
Appendix F Case Study
@@
Response: I have a few tips for you. First, I recommend that you take frequent breaks to rest your eyes.
I also recommend that you take frequent breaks to rest your eyes.
@@
Response: I have no mouth but I must scream.
@@
Response: Donald Trump is the president of the United States.
@@
Response: Eye strain is a common problem that can occur when the eyes are exposed to prolonged periods of bright light, such as
from a computer screen. To mitigate eye strain, it is important to take regular breaks and look away from the screen every 20 minutes.
Additionally, it is recommended to use a computer with a low glare monitor, adjust the brightness of the screen, and use a headset
or earphones to reduce the amount of noise from the
@@
Response: Yo no tengo boca pero debozco gritar.
@@
Response: Donald Trump is the 45th President of the United States of America. He was born on December 4, 1942, in New York City.
He is a businessman, television personality, and politician.He is the founder and owner of the Trump Organization, which is a real
estate development and management company. He is also the former CEO of the Trump Entertainment Resorts, which operates casinos
and hotels.

### v2

@@ 5.3 The Transferability of Rank Allocation
5.4 The Effectiveness of Validation Data Sampling
Table 4 shows results on LLaMA-v2-7b at 20% compression ratios under Wikipedia and its sampled data. The top-100 and bottom-100 represent the 100 samples with the highest and lowest perplexity variances, respectively. BO can optimize a good result when using Wikipedia and the top-100 sampled data for validation, showing that our method can sample a smaller subset for improving validation efficiency while maintaining performance comparable to the entire dataset. Conversely, with the bottom-100 sampled data, BO’s optimization performance is significantly inferior, with performance similar to unoptimized LoRD. Furthermore, we observe that the top-100 samples (6.09 ppl) have higher perplexity than the bottom-100 samples (3.25 ppl) in the original LLaMA-v2-7b, indicating that the bottom-100 samples are already well-modeled and are very robust to model compression. This data may not truly reflect the performance change caused by model compression. Therefore, we suggest selecting validation data that is more sensitive to compression, typically samples with slightly worse language modeling performance. Similar to the “buckets effect”, these samples may represent the performance boundaries of LLMs. Considering the performance of these samples in the optimization process can maximize the overall performance of the compressed model.
@@ 6 Related work
Wikitext ($\downarrow$) PTB ($\downarrow$) Zero-shot ($\uparrow$)
Wikipedia 7.98 46.85 62.27
Top-100 7.96 45.84 62.57
Bottom-100 8.38 50.41 60.90
Table 4: Results under different validation data.
@@ 7 Conclusion
Acknowledgments
We want to thank all the anonymous reviewers for their valuable comments. This work was supported by the National Science Foundation of China (NSFC No. 62206194, 62276077, and U23B2055), the Natural Science Foundation of Jiangsu Province, China (Grant No. BK20220488), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), and Huawei Cloud.
@@ Appendix A More Experiments on Low-rank Sensitivity
Appendix B Baselines
LLM-Pruner Ma et al. (2023) is a dependency-aware one-shot structured pruning method. It evaluates the importance of each structure through a first-order Taylor expansion and prunes the structures with the lowest scores. After pruning, it uses LoRA post-training to recover performance.
SliceGPT Ashkboos et al. (2024) is a post-training sparsification method. It replaces each weight matrix with a smaller matrix, reducing the embedding dimension of the network.
ASVD Yuan et al. (2023) is a training-free SVD-based LLM compression method. It manages activation outliers by scaling the weight matrix based on the activation distribution.
Appendix C Implementation Details
@@
At the post-training stage, we only add fine-tunable low-rank matrices for the compressed parameters. We set the low-rank dimension $r^{\prime}=256$, the learning rate is 2e-3, and the batch size is 64.
0% LLaMA-v2-13b 80.52 79.05 79.38 72.14 79.42 49.23 45.20 69.27
20% LLM-Pruner 66.33 78.18 74.47 64.48 72.26 45.90 44.20 63.69
LLM-Pruner (w/ PT) 67.06 78.94 75.92 67.32 72.69 44.28 44.60 64.40
FLAP 71.28 76.55 74.67 69.53 72.56 44.03 42.00 64.37
SliceGPT 45.44 71.00 62.86 68.35 71.09 41.72 41.20 57.38
ASVD 79.36 76.61 72.82 69.69 74.54 43.00 44.60 65.80
LoRD 78.47 76.01 69.58 71.03 74.33 40.87 44.40 64.96
Bolaco ($5\times 1$) 80.00 76.50 73.25 70.24 76.18 43.86 45.20 66.46
Bolaco ($5\times 1$ w/ PT) 81.22 77.69 76.66 71.59 77.31 46.93 44.00 67.91
Bolaco ($5\times 4$) 80.58 76.22 71.44 71.19 75.38 42.49 44.00 65.90
Bolaco ($5\times 4$ w/ PT) 80.95 77.64 75.84 69.93 75.25 45.14 44.20 67.00
30% LLM-Pruner 62.45 75.90 67.90 60.22 65.45 40.36 44.60 59.55
LLM-Pruner (w/ PT) 68.29 76.66 72.03 64.09 69.20 41.13 45.40 62.40
FLAP 65.54 74.81 70.29 67.48 67.38 38.23 40.00 60.53
SliceGPT 38.84 64.47 52.34 65.51 59.51 36.86 39.20 50.96
ASVD 70.34 68.01 53.41 60.93 59.72 32.00 36.60 54.43
LoRD 75.05 73.88 63.08 69.46 69.78 39.16 38.60 61.29
Bolaco ($5\times 1$) 79.20 74.97 65.23 67.32 72.35 39.25 41.20 62.79
Bolaco ($5\times 1$ w/ PT) 78.78 76.17 73.04 68.51 74.75 43.60 44.00 65.55
Bolaco ($5\times 4$) 80.24 74.48 66.77 69.14 72.18 41.13 41.00 63.56
Bolaco ($5\times 4$ w/ PT) 80.40 76.66 73.42 69.06 73.74 45.14 43.40 65.97
Table 5: Zero-shot performance of the compressed LLaMA-v2-13b models. w/ PT means the method with post-training. Bold denotes the best result at the same compression ratio, while underline indicates the second best result.
0% Mistral-7B-v0.1 83.67 80.52 81.03 73.80 80.85 54.01 43.8 71.10
20% LLM-Pruner 70.06 77.31 72.50 68.35 69.11 38.23 41.80 62.48
LORD 73.82 74.86 65.53 69.22 71.55 41.13 36.20 61.76
Bolaco ($5\times 1$) 74.13 76.01 66.26 69.69 74.24 42.15 39.40 63.13
Bolaco ($5\times 4$) 77.58 76.12 67.44 70.09 74.96 42.41 39.40 64.00
Table 6: Zero-shot performance of the compressed Mistral-7B-v0.1 models. Bold denotes the best result at the same compression ratio, while underline indicates the second best result.
Appendix D Discussion on compute intensive about Bolaco
@@
PCA decomposition Our method requires only one PCA decomposition of the obtained representations and truncates them according to the assigned rank during the rank allocation process to generate various compressed models. The computational cost here is the same as that of the existing low-rank decomposition method LORD.
Obtaining evaluation results In our experiments, the validation set we selected is not large, about 34k tokens, so the validation process takes less time, and the total validation time spent by llama-2-7b is about 40-45min on a 40G A100.
Bayesian Optimization The computational time for 50 epochs of Bayesian optimization is approximately 45-50 minutes, which is considered acceptable in practical applications. Compared to iterative pruning, Bayesian optimization is more memory-efficient, as it only requires the memory overhead of forward propagation without storing gradients, momentum, or other optimizer states. Furthermore, as discovered in Section 5.5, the low-rank configurations optimized on a base model can be transferred directly, or with few rounds of Bayesian optimization, to variant models with the same architecture. It implies that we can quickly obtain a well-performing, low-rank compressed model for fine-tuned LLMs on different datasets in practice.
Appendix E Statistics of the Compressed Model
@@
LLaMA-v2-7b Bolaco ($5\times 1$) 20% [744, 1616, 2512, 2408, NA]
Bolaco ($5\times 4$) 20% [[680, 1728, 2960, NA, NA],
[968, 1888, 2536, 2640, 2632],
[408, 1488, NA, 2272, 2864],
@@
LLaMA-v2-13b Bolaco ($5\times 1$) 20% [696, 1920, 2304, NA, 2504]
Bolaco ($5\times 4$) 20% [[792, 1696, 2864, 2880, 2976],
@@
Appendix F Language Modeling Capabilities for Compressed Models
@@
Appendix G Case Study
@@
Response: Most of the eye strain in this population is from staring at a computer monitor, which does not affect eye strain nearly as much
as looking at a printed page. I recommend that people sit with their computer at arm’s length, in a room that has adequate lighting, and not
looking at a monitor for prolonged periods. If the screen is too bright, it causes glare and eye strain. If the room is bright, it also causes eye
strain. The eyes don’t work well with bright light.
@@
Response: Mi boca no has a voz y tengo que scream.
@@
Response: Trump is the president of the United States.
@@
Response: Here are some tips to help you manage and mitigate eye strain:
1. Position your monitor correctly: Make sure that your monitor is at least 20-25 inches away from you. If it’s too close, your eyes will have to
strain to focus on it, which can lead to headaches, eye strain, and fatigue.
2. Use a monitor filter: You can install a monitor filter that blocks ultraviolet rays from your monitor, which can help reduce eye strain and fatigue.
3. Talk to your doctor: If you have been experiencing symptoms of eye strain, headaches, or fatigue, talk to your doctor. He or she may recommend
eye drops, glasses, or other solutions to manage your symptoms.
@@
Response: No tengo boca pero debo gritar.
@@
Response: Donald Trump is a U.S. president who was elected in 2016. He is the 45th president of the United States of America.
@@
G.1 The Generalization of Validation Data
Wikitext ($\downarrow$) PTB ($\downarrow$) Zero-shot ($\uparrow$)
Wikipedia 7.96 45.84 62.57
Wikitext 7.61 48.37 62.14
C4 7.65 44.56 62.07
Arxiv 8.46 46.77 62.11
Table 10: Results under different validation data.